Can you help me with this equation?
A rectangle has a perimeter of 38 feet. The length is 1 foot less than 3 times the width. Find the dimensions of the rectangle.
Thank You

Answer by checkley77(12844)

2L+2W=38
L=3W-1
2(3W-1)+2W=38
6W-2+2W=38
8W=38+2
8W=40
W=40/8
W=5 ANS. FOR THE WIDTH.
L=3*5-1
L=15-1
L=14 ANS. FOR TH LENGTH.
PROOF:
2*14+2*5=38
28+10=38
38=38